Can you show how to get answer to this word problem please
An aircraft 200 kilometers away is flying toward an air base at 350 kilometers per hour. How long will it take a plane flying from the base at 450 kilometers per hour to intercept the first aircraft?

Answer by checkley77(12844)

D=RT
200=(350+450)T
200=800T
T=200/800
T=2/8=1/4 THEY WILL MEET IN 15 MINUTES.